1-PHENYLPYRROLE A TICT MOLECULE?.^{a}

\underline{JESSICA A. THOMAS}, JUSTIN W. YOUNG, ADAM J. FLEISHER, CASEY L. CLEMENTS, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260; LEONARDO ALVAREZ-VALTIERRA, Division de Ciencias e Ingenierias, Campus Leon, Universidad de Guanajuato, Leon, Gto. 37150, Mexico.

1-Phenylpyrrole has been identified by several authors as a molecule that displays twisted intramolecular charge-transfer (TICT) character in the condensed phase. However, Suzuka and coworkers ^{b}, using single vibronic level dispersed fluorescence, observed no evidence of TICT activity in the gas phase. In this work, 1-phenylpyrrole is studied by the high resolution methods previously applied to DMABN ^{c} to determine the extent of its motion along different possible TICT coordinates when it absorbs light.

aWork supported by NSF CHE-0615755

bK. Okuyama, Y. Numata, S. Odawara, and I. Suzuka, J. Chem. Phys., 109, 7185 (1998) and references therein.

cA. E. Nikolaev, G. Myszkiewwicz, G. Berden, L. Meerts, J. F. Pfanstiel, and D. W. Pratt, J. Chem. Phys., 122, 084309 (2005).